# LP-ICP: General Localizability-Aware Point Cloud Registration for Robust Localization in Extreme Unstructured Environments

### v2

@@ Abstract
The Iterative Closest Point (ICP) algorithm is a crucial component of LiDAR-based SLAM algorithms. However, its performance can be negatively affected in unstructured environments that lack features and geometric structures, leading to low accuracy and poor robustness in localization and mapping. It is known that degeneracy caused by the lack of geometric constraints can lead to errors in 6-DOF pose estimation along ill-conditioned directions. Therefore, there is a need for a broader and more fine-grained degeneracy detection and handling method. This paper proposes a new point cloud registration framework, LP-ICP, that combines point-to-line and point-to-plane distance metrics in the ICP algorithm, with localizability detection and handling. LP-ICP consists of a localizability detection module and an optimization module. The localizability detection module performs localizability analysis by utilizing the correspondences between edge points (with low local smoothness) to lines and planar points (with high local smoothness) to planes between the scan and the map. The localizability contribution of individual correspondence constraints can be applied to a broader range. The optimization module adds additional soft and hard constraints to the optimization equations based on the localizability category. This allows the pose to be constrained along ill-conditioned directions, with updates either tending towards the constraint value or leaving the initial estimate unchanged. This improves accuracy and reduces fluctuations. The proposed method is extensively evaluated through experiments on both simulation and real-world datasets, demonstrating higher or comparable accuracy than the state-of-the-art methods. The dataset and code of this paper will also be open-sourced at https://github.com/xuqingyuan2000/LP-ICP.
@@
This paper was motivated by addressing the challenges of Simultaneous Localization and Mapping (SLAM) that use LiDAR as one of the sensors in extreme unstructured environments such as planetary-like environments and underground tunnels. Due to the lack of features and geometric structures in these environments, the performance of pose estimation based on the Iterative Closest Point (ICP) algorithm is limited in the degenerate directions. This leads to low accuracy and poor robustness in localization and mapping. Most existing degeneracy detection methods have limited applicability, such as those that only work with point-to-plane ICP algorithms, or are not fine-grained. This paper proposes a new point cloud registration algorithm that analyzes the localizability of point-to-line and point-to-plane correspondences to detect and handle degeneracy in pose estimation. The proposed localizability detection method has the potential to be extended to other variants of ICP or multi-sensor fusion frameworks. The optimization module incorporates both soft and hard constraints to improve accuracy and reduce fluctuations. We have demonstrated through extensive experiments that the proposed algorithm performs accurate localizability detection and robust pose estimation in challenging environments, achieving higher or comparable accuracy in localization and mapping.
@@ I Introduction
Point cloud registration is a crucial step in LiDAR-based SLAM approaches. It aligns two sets of point clouds by finding the optimal rotation and translation [4, 5]. The Iterative Closest Point (ICP) algorithm is commonly used for scan-to-scan or scan-to-map registration to estimate the robot’s pose. Common distance metrics include point-to-point [6], point-to-line [7], and point-to-plane [8], [9]. These methods have been widely adopted in many advanced SLAM systems [2, 10, 11, 12, 13]. In recent years, the ICP algorithm has continued to undergo further research [14, 15, 16].
In typical urban or indoor environments, there are generally sufficient geometric constraints for existing point cloud registration methods. Therefore, the localization accuracy in these environments is quite satisfactory. However, extreme unstructured environments, such as planetary-like environments [17, 18] on Mars or the Moon, as well as underground tunnels [19], as shown in Fig. 1, are characterized by self-similarity and sparse geometrical structures. This results in insufficient constraints being provided for ICP optimization. The optimization is rendered degenerate in certain directions of the 6-DoF pose, leading to decrease in the accuracy of localization and mapping.
@@
Different geometric constraint information, such as planes in various orientations or their edges, has varying impacts on the solution of ICP optimization. The evaluation of such impacts and the handling of different levels of degeneracy should be incorporated into the framework to improve the nature of the solution. Zhang et al. [20] were the first to introduce the degeneracy factor into the optimization problem of state estimation, which can be used to detect the overall degeneracy level in each direction of the 6-DoF pose. However, it can only assess the overall degeneracy in a given direction and cannot be used to calculate the localizability contribution of individual constraints within it. Additionally, the binary judgment of degeneracy relies on threshold settings, which may lead to insufficient utilization of the constraints information. Tuna et al. [21] recently introduced localizability detection of individual constraints in point-to-plane ICP for point cloud registration and established a ternary judgment for degeneracy, referred to as X-ICP. However, their method does not utilize point-to-line localizability information, which limits the potential for further accuracy improvement. Moreover, the formula for calculating the localizability contribution of individual constraints cannot be generalized to broader ICP algorithms and their variants. Additionally, X-ICP introduces hard constraints in partially localizable optimization directions. This increases the risk of fluctuations or even significant drift in the estimated trajectory in some SLAM systems, potentially leading to localization failure.
@@
The Optimization-Module first calculates additional constraints, utilizing localizability information from geometric constraints and correspondences with high localizability contributions. Then, based on the localizability categories from the LocDetect-Module, soft and hard constraints are added to the optimization equation. This ensures that the pose in non-localizable directions is not updated, maintaining the initial estimate unchanged. In partially localizable directions, the pose is updated under additional constraints, tending toward the constraint value rather than being strictly equal to it. This reduces the risk of large pose fluctuations and drift, which could lead to localization failure. This is one of the differences between our approach and [21]. For fully localizable directions, the pose update remains unaffected. The aim is to effectively utilize more geometric constraint information while attempting to ensure relatively stable pose updates.
@@
The proposed framework has been tested in multiple experiments on challenging simulation and real-world datasets. The experiments suggest that LP-ICP performs well in extreme unstructured environments and can improve the robot’s localization capability and mapping accuracy in degraded environments. Compared to the state-of-the-art approaches [20], [21], the proposed framework achieves higher or comparable accuracy in all experiments.
@@
  • We propose a new ICP algorithm framework with localizability detection and handling, utilizing point-to-line and point-to-plane localizability information. It is used to detect and handle degeneracy in extreme unstructured environments.
@@
  • Comparing with state-of-the-art methods, a variety of experiments verify the efficacy of the proposed framework.
@@ II-A ICP Algorithm and Its Variants
Point cloud registration aligns two or more sets of point clouds into a unified coordinate system by solving for the pose transformation between them. ICP is a commonly used method for point cloud registration. The ICP algorithm iteratively solves for the pose transformation between the source and target point clouds by minimizing a cost function, obtaining the optimal solution. Over the years, the ICP algorithm has evolved to include cost functions with different metrics, such as point-to-point [6], point-to-line [7], and point-to-plane [8], [9]. Segal et al. [22] combined point-to-point and point-to-plane ICP into a probabilistic framework. Furthermore, Billings et al. [23] introduced a noise model for further extension. Serafin and Grisetti [24] proposed NICP, which incorporates angular errors of normal vectors into the cost function. Li et al. [25] considered both the angles and distances of local normals while introducing a point-to-surface version of ICP. Additionally, there are learning-based methods, such as [26]. Currently, the ICP algorithm combining point-to-line and point-to-plane distance metrics remains one of the preferred choices. It has been applied in many advanced SLAM systems, such as [2, 10, 11, 27].
@@ II-B Degeneracy Detection and Handling
Zhang et al. [20] pioneered the concept of the degeneracy factor, using the eigenvalues of the Hessian matrix to evaluate the overall degeneracy in the six directions of pose estimation. By solution remapping, they ensure that poses in degenerate directions do not update, thereby mitigating the negative impact of degeneracy. This method has been adopted in several LiDAR-based SLAM frameworks, such as [2, 11, 30, 31, 32, 33]. Hinduja et al. [34] used the condition number to set a threshold for determining degeneracy in the point-to-plane ICP algorithm. During optimization, they applied the solution remapping technique. The results were incorporated into the SLAM pose graph optimization as partial loop closure factors, ensuring that ICP and pose graph updates and optimization are performed only in well-constrained directions of the state space. Lee et al. [35], building on [20], utilized a chi-square distribution to non-heuristically set thresholds for the three most degenerate directions in LiDAR odometry. Wen et al. [36] set a threshold for binary detection of degeneracy by using the total residuals of the ICP algorithm in LiDAR odometry and the pose estimation results of the Visual-Inertial Odometry (VIO). In the degenerate direction of the LiDAR odometry, only visual measurements and IMU residuals are used for estimation. However, these methods cannot be used to evaluate the localizability contribution of individual constraints. Furthermore, the binary degeneracy detection depends on threshold settings, which may result in insufficient utilization of constraint information.
@@
Zhen et al. [37] proposed a method that constructs an information matrix using the plane normal vectors from an existing map. They then use its eigenvalues to assess the localizability in the translational directions. However, they did not provide a detailed explanation or handling approach for the metrics. Furthermore, Zhen and Scherer [38] in the point-to-plane ICP algorithm used the sensitivity of LiDAR measurements w.r.t. the 6-DoF pose as a measure of localizability to assess the strength of the constraints. They also utilized the geometric information from the point and plane normals. To handle low localizability, they chose to incorporate measurements from other sensors for compensation. Tuna et al. [21] recently introduced localizability detection into the point-to-plane ICP algorithm, utilizing point and plane normal information, referred to as X-ICP. They pioneered a ternary degeneracy judgment mechanism for each direction of the pose solution, thereby incorporating more constraint information. The localizability calculation formula for individual constraints is explained from a geometric or force perspective. However, further mathematical derivation and explanation from the perspective of degeneracy are necessary. Furthermore, in some SLAM systems, X-ICP optimization uses hard constraints in partially localizable directions, which increases the risk of trajectory fluctuations or even significant drift, potentially leading to localization failure. Hatleskog and Alexis [39] proposed a method in the point-to-plane ICP algorithm that considers the noise in both point and plane normals. By accounting for the probabilistic characteristics of the noise entering the Hessian, their approach helps to detect degeneracy and reduces the impact of noise in degenerate directions. These studies provide more fine-grained detection of localizability and enable the evaluation of the localizability contribution of individual constraints. However, they are limited to the point-to-plane ICP algorithm and do not utilize the localizability information from point-to-line correspondences, which limits further improvements in accuracy.
@@
Nubert et al. [40] proposed using a learning-based approach to detect the localizability of LiDAR. The method is trained with simulated data and utilizes the point-to-plane ICP algorithm. Degeneracy detection does not require heuristic threshold adjustments, but it still relies on a binary decision. Moreover, it is only applicable to scan-to-scan point cloud registration. There are also learning-based methods such as [41, 42]. However, their reliance on labeled ground truth data and high computational demands limits their use in resource-constrained systems.
@@ III-A Optimization-Based Point Cloud Registration
An optimization-based state estimation problem can be formulated by solving a function:
@@
$f_{pj}(\boldsymbol{x})=\|(\boldsymbol{R}\boldsymbol{p}_{j}^{L}+\boldsymbol{t}-\boldsymbol{q}_{j}^{M})\cdot\boldsymbol{n}_{j}^{M}\|.$ (5)
@@
Specifically, the alignment of the current scan point cloud $\boldsymbol{p}^{L}$ in the LiDAR frame (denoted as $L$) with the map point cloud $\boldsymbol{q}^{M}$ in the map frame (denoted as $M$) is achieved through the transformation $\boldsymbol{T}=\begin{bmatrix}\boldsymbol{R}\;|\;\boldsymbol{t}\end{bmatrix}$ (or $\boldsymbol{x}=(\boldsymbol{r}^{T}\;\boldsymbol{t}^{T})^{T}\in\mathbb{R}^{6}$). Here,$\boldsymbol{R}\in SO(3)$ denote the rotation matrix and $\boldsymbol{t}\in\mathbb{R}^{3}$ denote the translation vector. $\boldsymbol{p}_{i}^{L}$ and $\boldsymbol{p}_{j}^{L}$ are the edge and planar points extracted from the LiDAR frame, respectively. $\boldsymbol{q}_{i}^{M}$ and $\boldsymbol{q}_{j}^{M}$ are the corresponding matching points in the map frame, typically found using a search method, such as k-d tree search. $\boldsymbol{l}_{j}^{M}$ is the unit direction vector of the edge line passing through $\boldsymbol{q}_{i}^{M}$, and $\boldsymbol{n}_{j}^{M}$ is the unit normal vector of the plane passing through $\boldsymbol{q}_{j}^{M}$.
@@ III-B Degeneracy Detection
Here, $\lambda$ is the eigenvalue of the matrix $\boldsymbol{A}^{T}\boldsymbol{A}$, where $\boldsymbol{A}$ comes from the following equation (i.e. the linearized form of (1)):
$\arg\min_{\boldsymbol{x}}\|\boldsymbol{A}\boldsymbol{x}-\boldsymbol{b}\|.$ (7)
The Gauss-Newton method is a commonly used approach for solving optimization problems. In optimization problems solved using the Gauss-Newton method, $\lambda$ is the eigenvalues corresponding to the Hessian matrix of the original optimization problem. $\lambda$ can be used to evaluate whether optimization in a specific direction suffers from degeneracy. However, it cannot assess the localizability contribution of individual correspondence constraints in that direction. As a result, geometric information cannot be more effectively utilized in the optimization process to improve localization accuracy. Therefore, obtaining more fine-grained localizability detection results is one of the objectives of this study.

### v3

@@ Abstract
The Iterative Closest Point (ICP) algorithm is a crucial component of LiDAR-based SLAM algorithms. However, its performance can be negatively affected in unstructured environments that lack features and geometric structures, leading to low accuracy and poor robustness in localization and mapping. It is known that degeneracy caused by the lack of geometric constraints can lead to errors in 6-DOF pose estimation along ill-conditioned directions. Therefore, there is a need for a broader and more fine-grained degeneracy detection and handling method. This paper proposes a new point cloud registration framework, LP-ICP, that combines point-to-line and point-to-plane distance metrics in the ICP algorithm, with localizability detection and handling. Rather than relying solely on point-to-plane localizability information, LP-ICP enhances the localizability analysis by incorporating a point-to-line metric, thereby exploiting richer geometric constraints. It consists of a localizability detection module and an optimization module. The localizability detection module performs localizability analysis by utilizing the correspondences between edge points (with low local smoothness) to lines and planar points (with high local smoothness) to planes between the scan and the map. The localizability contribution of individual correspondence constraints can be applied to a broader range. The optimization module adds additional soft and hard constraints to the optimization equations based on the localizability category. This allows the pose to be constrained along ill-conditioned directions. The proposed method is evaluated on simulation and real-world datasets, showing comparable or better accuracy than the state-of-the-art methods in tested scenarios. Observed variations in partially localizable directions suggest the need for further investigation on robustness and generalizability.
@@
This paper was motivated by addressing the challenges of Simultaneous Localization and Mapping (SLAM) that use LiDAR as one of the sensors in extreme unstructured environments such as planetary-like environments and underground tunnels. Due to the lack of features and geometric structures in these environments, the performance of pose estimation based on the Iterative Closest Point (ICP) algorithm is limited in the degenerate directions. This leads to low accuracy and poor robustness in localization and mapping. Most existing degeneracy detection methods have limited applicability or are not fine-grained. This paper proposes a new point cloud registration algorithm that additionally analyzes the localizability of point-to-line correspondences to detect and handle degeneracy in pose estimation. The proposed localizability detection method has the potential to be extended to other variants of ICP or multi-sensor fusion frameworks. The optimization module incorporates both soft and hard constraints based on the localizability analysis. Extensive experiments demonstrate that the proposed algorithm effectively detects localizability and achieves better or comparable performance to the state-of-the-art methods in challenging environments.
@@ I Introduction
Point cloud registration is a crucial step in LiDAR-based SLAM approaches. It aligns two sets of point clouds by estimating the rotation and translation between them [4, 5]. The Iterative Closest Point (ICP) algorithm is commonly used for scan-to-scan or scan-to-map registration to estimate the robot’s pose. Common distance metrics include point-to-point [6], point-to-line [7], and point-to-plane [8], [9]. These methods have been widely adopted in many advanced SLAM systems [2, 10, 11, 12, 13]. In recent years, the ICP algorithm has continued to undergo further research [14, 15, 16].
In geometrically rich environments, such as urban streets or campus classrooms, there are generally sufficient geometric constraints for existing point cloud registration methods. Therefore, the localization accuracy in these environments is quite satisfactory. However, extreme unstructured environments, such as planetary-like environments [17, 18] on Mars or the Moon, and underground tunnels [19, 20], as shown in Fig. 1, as well as vast open areas, are characterized by self-similarity and sparse geometrical structures. This results in insufficient constraints being provided for ICP optimization. The optimization is rendered degenerate in certain directions of the 6-DoF pose, leading to decrease in the accuracy of localization and mapping.
@@
Different geometric constraint information, such as planes in various orientations or their edges, has varying impacts on the solution of ICP optimization. The evaluation of such impacts and the handling of different levels of degeneracy should be incorporated into the framework to improve the nature of the solution. Zhang et al. [21] were the first to introduce the degeneracy factor into the optimization problem of state estimation, which can be used to detect the overall degeneracy level in each direction of the 6-DoF pose. They also employed the solution remapping after the binary judgment of degeneracy. However, it can only assess the overall degeneracy in a given direction and cannot be used to calculate the localizability contribution of individual constraints within it. Tuna et al. [22] recently introduced localizability detection of individual constraints in point-to-plane ICP for point cloud registration and established a ternary judgment for degeneracy, referred to as X-ICP. However, their method does not utilize point-to-line localizability information, which limits the potential for further accuracy improvement. Moreover, the formula for calculating the localizability contribution of individual constraints cannot be generalized to broader ICP algorithms and their variants.
@@
The Optimization-Module first calculates additional constraints, utilizing localizability information from geometric constraints and correspondences with high localizability contributions. Then, based on the localizability categories from the LocDetect-Module, soft and hard constraints are added to the optimization equation. This ensures that the pose in non-localizable directions is not updated, maintaining the initial estimate unchanged. In partially localizable directions, the pose is updated under additional constraints. For fully localizable directions, the pose update remains unaffected.
@@
The proposed framework has been tested in multiple experiments on challenging simulation and real-world datasets. The experiments suggest that LP-ICP performs well in extreme unstructured environments and shows potential to improve the robot’s localization capability and mapping accuracy in degraded environments. Compared to the state-of-the-art approaches [21], [22], the proposed framework achieves comparable or improved accuracy in our tested scenarios.
@@
  • We propose a new ICP algorithm framework with localizability detection and handling, utilizing point-to-line and point-to-plane localizability information. It introduces a novel point-to-line localizability analysis to help detect and handle degeneracy in extreme unstructured environments.
@@
  • A variety of experiments demonstrate the effectiveness of the proposed framework under our experimental conditions, in comparison with state-of-the-art methods.
@@ II-A ICP Algorithm and Its Variants
Point cloud registration aligns two or more sets of point clouds into a unified coordinate system by solving for the pose transformation between them. ICP is a commonly used method for point cloud registration. The ICP algorithm iteratively solves for the pose transformation between the source and target point clouds by minimizing a cost function, obtaining the optimal solution. Over the years, the ICP algorithm has evolved to include cost functions with different metrics, such as point-to-point [6], point-to-line [7], and point-to-plane [8], [9]. Segal et al. [23] combined point-to-point and point-to-plane ICP into a probabilistic framework. Furthermore, Billings et al. [24] introduced a noise model for further extension. Appropriately combining different error metrics can enhance the robustness of the algorithm across varying environments [25]. Lee et al. proposed GenZ-ICP [25], which adaptively sets the weights of point-to-point and point-to-plane error metrics in the optimization problem based on the ratio of planar points to non-planar points. This mitigates the negative impact of optimization degeneracy in point-to-plane ICP in degraded environments. Serafin and Grisetti [26] proposed NICP, which incorporates angular errors of normal vectors into the cost function. Li et al. [27] considered both the angles and distances of local normals while introducing a point-to-surface version of ICP. Additionally, there are learning-based methods, such as [28]. Currently, the ICP algorithm combining point-to-line and point-to-plane distance metrics by utilizing edge and planar features in the environment remains one of the preferred choices. It has been applied in many advanced SLAM systems, such as [2, 10, 11, 29].
@@ II-B Degeneracy Detection and Handling
Zhang et al. [21] pioneered the concept of the degeneracy factor, using the eigenvalues of the Hessian matrix to evaluate the overall degeneracy in the six directions of pose estimation. By solution remapping, they ensure that poses in degenerate directions do not update, thereby mitigating the negative impact of degeneracy. This method has been adopted in several LiDAR-based SLAM frameworks, such as [2, 11, 32, 33, 34, 35]. Hinduja et al. [36] used the condition number to set a threshold for determining degeneracy in the point-to-plane ICP algorithm. During optimization, they applied the solution remapping technique. The results were incorporated into the SLAM pose graph optimization as partial loop closure factors, ensuring that ICP and pose graph updates and optimization are performed only in well-constrained directions of the state space. Lee et al. [37], building on [21], utilized a chi-square distribution to non-heuristically set thresholds for the three most degenerate directions in LiDAR odometry. McDermott and Rife [38] proposed ICET, which enhances the Normal Distributions Transform (NDT) by introducing a condition number framework to detect global degeneracy and suppress solutions along degenerate directions while preserving information in useful directions. This helps reduce geometric ambiguity along large flat surfaces. Wen et al. [39] set a threshold for binary detection of degeneracy by using the total residuals of the ICP algorithm in LiDAR odometry and the pose estimation results of the Visual-Inertial Odometry (VIO). In the degenerate direction of the LiDAR odometry, only visual measurements and IMU residuals are used for estimation. However, these methods are not fine-grained enough and cannot be used to evaluate the localizability contribution of individual constraints.
@@
Zhen et al. [40] proposed a method that constructs an information matrix using the plane normal vectors from an existing map. They then use its eigenvalues to assess the localizability in the translational directions. However, they did not provide a detailed explanation or handling approach for the metrics. Furthermore, Zhen and Scherer [41] in the point-to-plane ICP algorithm used the sensitivity of LiDAR measurements w.r.t. the 6-DoF pose as a measure of localizability to assess the strength of the constraints. They also utilized the geometric information from the point and plane normals. To handle low localizability, they chose to incorporate measurements from other sensors for compensation. Tuna et al. [22] recently introduced localizability detection into the point-to-plane ICP algorithm, utilizing point and plane normal information, referred to as X-ICP. They pioneered a ternary degeneracy judgment mechanism for each direction of the pose solution, thereby incorporating more constraint information. The localizability calculation formula for individual constraints is explained from a geometric or force perspective. However, further mathematical derivation and explanation from the perspective of degeneracy are necessary. Chen et al. [42] adopted the degeneracy detection method proposed in [22] for LiDAR-inertial odometry and incorporated the computed additional constraints into the ESIKF update. Tuna et al. [43] proposed and introduced new methods such as inequality constraints, truncated SVD, and Tikhonov regularization into the field of degeneracy-aware point cloud registration. They also evaluated and analyzed the effectiveness of these different constraint types in comparison with existing methods. Hatleskog and Alexis [44] proposed a method in the point-to-plane ICP algorithm that considers the noise in both point and plane normals. By accounting for the probabilistic characteristics of the noise entering the Hessian, their approach helps to detect degeneracy and reduces the impact of noise in degenerate directions. These studies provide more fine-grained detection of localizability and enable the evaluation of the localizability contribution of individual constraints. However, they are limited to the point-to-plane ICP algorithm and do not utilize the localizability information from point-to-line correspondences, which limits further improvements in accuracy.
@@
Nubert et al. [45] proposed using a learning-based approach to detect the localizability of LiDAR. The method is trained with simulated data and utilizes the point-to-plane ICP algorithm. Degeneracy detection does not require heuristic threshold adjustments, but it is only applicable to scan-to-scan point cloud registration. There are also learning-based methods such as [46, 47]. However, their reliance on labeled ground truth data and high computational demands limits their use in resource-constrained systems.
@@ III-A Optimization-Based Point Cloud Registration
Nonlinear optimization plays a crucial role in many SLAM systems[1, 48, 2]. An optimization-based state estimation problem can be formulated by solving a function:
@@
$f_{pj}(\boldsymbol{x})=(\boldsymbol{R}\boldsymbol{p}_{j}^{L}+\boldsymbol{t}-\boldsymbol{q}_{j}^{M})\cdot\boldsymbol{n}_{j}^{M}.$ (5)
@@
Specifically, the alignment of the current scan point cloud $\boldsymbol{p}^{L}$ in the LiDAR frame (denoted as $L$) with the map point cloud $\boldsymbol{q}^{M}$ in the map frame (denoted as $M$) is achieved through the transformation $\boldsymbol{T}=\begin{bmatrix}\boldsymbol{R}\;|\;\boldsymbol{t}\end{bmatrix}$ (or $\boldsymbol{x}=(\boldsymbol{r}^{T}\;\boldsymbol{t}^{T})^{T}\in\mathbb{R}^{6}$). Here, $\boldsymbol{R}\in SO(3)$ denotes the rotation matrix and $\boldsymbol{t}\in\mathbb{R}^{3}$ denotes the translation vector. $\boldsymbol{p}_{i}^{L}$ and $\boldsymbol{p}_{j}^{L}$ are the edge and planar points extracted from the LiDAR frame, respectively. $\boldsymbol{q}_{i}^{M}$ and $\boldsymbol{q}_{j}^{M}$ are the corresponding matching points in the map frame, typically found using a search method, such as k-d tree search. $\boldsymbol{l}_{j}^{M}$ is the unit direction vector of the edge line passing through $\boldsymbol{q}_{i}^{M}$, and $\boldsymbol{n}_{j}^{M}$ is the unit normal vector of the plane passing through $\boldsymbol{q}_{j}^{M}$.
@@ III-B Degeneracy Detection
Here, $\lambda$ is the eigenvalue of the matrix $\boldsymbol{A}^{T}\boldsymbol{A}$, which can be computed using eigenvalue decomposition. Note that the eigenvalues of $\boldsymbol{A}^{T}\boldsymbol{A}$ are typically referred to as singular values and computed using the singular value decomposition. $\boldsymbol{A}$ comes from the following equation (i.e. the linearized form of (1)):
$\arg\min_{\boldsymbol{x}}\|\boldsymbol{A}\boldsymbol{x}-\boldsymbol{b}\|.$
The Gauss-Newton method is a commonly used approach for solving optimization problems. In optimization problems solved using the Gauss-Newton method, $\lambda$ is the eigenvalues corresponding to the Hessian matrix of the original optimization problem. $\lambda$ can be used to evaluate whether optimization in a specific direction suffers from degeneracy. However, it cannot assess the localizability contribution of individual correspondence constraints in that direction. Therefore, obtaining more fine-grained localizability detection results is one of the objectives of this study.